Sqrt (3+2) + Sqrt (2x-1)= 5

Answer by ewatrrr(24785)

Sqrt (3+2) + Sqrt (2x-1)= 5
Sqrt (2x-1)= 5 - Sqrt (5)
Sqrt (2x-1) = 2.764 (nearest thousandth)
2x - 1 = 7.639
x = 8.639/2